OpenStudy (anonymous):

What is the equation of a line passing through the point (a,b) and having a slope of b?

OpenStudy (amriju):

Take the standard eqn of the line thats y=mx+c..m is slope...so (a,b) must satisfy the eqn...and mis provided to be b...so b=ab+c...since c is a constant...we have c=b(1-a)...so the eqn is y=bx+b(1-a)..

OpenStudy (anonymous):

use the point slope formula \[y-y_1=m(x-x_1)\] with \(m=b, x_1=a, y_1=b\)

OpenStudy (anonymous):

you get \[y-b=b(x-a)\]

OpenStudy (anonymous):

Wait that's not one of the answers. I have.. Y= -bx - b (1 - a) Y= bx - b(a-1) Y= 1/b x + b(a-1) Y= -1/box + b(1-a)

OpenStudy (anonymous):

so you can see that each of your answers has \(y=\text{something}\) your job is to take \[y-b=b(x-a)\] and solve for \(y\)

OpenStudy (amriju):

b dear...b(1-a)=-b(a-1) isn't it??

OpenStudy (anonymous):

actually you don't have to do anything your are told that the slope is \(b\) only one of your possible answers has slope \(b\)

OpenStudy (anonymous):

\[y= -bx - b (1 - a)\\ y= bx - b(a-1)\\ y= 1/b x + b(a-1) \\y= -1/bx + b(1-a)\]

OpenStudy (amriju):

lololol..:P...cool..

OpenStudy (anonymous):

only the second one has slope \(b\) so either that is the right choice or none of them are